OpenStudy (anonymous):

In a random sample of 63 women at a company, the mean salary is $48,902 with a standard deviation of $5270. In a random sample of 50 men at the company, the mean salary is $53,454 with a standard deviation of $4677. Construct a 95% confidence interval for the difference between the mean salaries of all women and men at the company. (Points : 4) ($2134.05, $6969.95) ($2715.11, $6388.89) ($3614.81, $5489.19) ($4083.40, $5020.60)

OpenStudy (kropot72):

The 95% confidence interval for the difference between the mean salaries is found from the following: \[\large (\bar{x}_{1}-\bar{x}_{2})\pm1.96\sqrt{\frac{s _{1}^{2}}{n _{1}}+\frac{s _{2}^{2}}{n _{2}}}\] where s1 is the standard deviation of the first sample and s2 is the standard deviation of the second sample.

OpenStudy (kropot72):

@Kyliejones Just substitute the given values and you will find the result agrees with one of the answer choices.

OpenStudy (perl):

*